Juana bought 16 used paperback books for $10.95. Some cost $.60 and the rest cost $.75 each. How many of each did she buy?
Answer by stanbon(75887)

Juana bought 16 used paperback books for $10.95. Some cost $.60 and the rest cost $.75 each. How many of each did she buy?
----
Equations:
a + b = 16
0.6a + 0.75b = 10.95
------
a + b = 16
60a + 75b = 1095
----------
a = 16-b
12a + 15b = 219
---
Substitute for "a" and solve for "b":
12(16-b) + 15b = 219
192 - 12b + 15b = 219
3b = 27
b = 9 (# of 0.60 books)
----
a = 16-9 = 7 (# of 0.75 books)
===================================
Cheers,
Stan H.
=============